an object travels in a circular orbit. If the speed of the object is halved, it's centripetal acceleration will be...


1
Expert's answer
2022-03-11T09:00:42-0500

We know that

Centripetal acceleration

"a_c=\frac{v^2}{r}"


"v'=\frac{v}{2}"

New centripetal acceleration

"a_c'=\frac{v'^2}{r}"

"a'_c=\frac{v^2}{4r}"

"a'_c=\frac{a_c}{4}"